Find the real solutions of the equation by factoring: x^3+9x^2-x-9=0
Answer by nerdybill(7384)

x^3+9x^2-x-9=0
group terms:
(x^3+9x^2) - (x+9) = 0
x^2(x+9) - (x+9) = 0
(x+9)(x^2-1) = 0
(x+9)(x^2-1^2) = 0
(x+9)(x+1)(x-1) = 0
x = {-9,-1,1}